# Modeling wildfire dynamics through a physics-based approach incorporating fuel moisture and landscape heterogeneity

### v1

@@ Acknowledgments
This work has been funded by the Spanish Ministry of Science, Innovation and Universities - Agencia Estatal de Investigación (10.13039/501100011033) and FEDER-EU under project-nr. PID2022-141051NA-I00. This work was partially funded by the Aragón Government, DGA (Spain), through the Fondo Europeo de Desarrollo Regional (FEDER) under project-nr. T32-23R. I. Özgen-Xian’s work is funded by the Tenure-Track-Programm by the German Federal Ministry of Education and Research.
@@ Appendix A Energy conservation equations in the model
In this section, we present some of the equations for the conservation of energy and the assumptions made to develop the model herein presented. Refer to [8] for a complete description of these equations. For the sake of clarity, notation has been simplified and subscripts for the material properties for the fluid and solid cases have been omitted when analyzed independently.
@@
Now, by expanding the left-hand side of Eq. (86) and making use of the mass conservation equation—i.e. $\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{v})=0$, we can rewrite Eq. (86) in non-conservative form
@@
By expressing enthalpy differential in terms of pressure and temperature [8], we can rewrite Eq. (87) as
@@
where $c_{p}$ is the specific heat at constant pressure. If we consider an ideal gas, we can notice the relation $\left(\frac{\partial\ln\rho}{\partial\ln T}\right)_{p}=-1$ and the previous equation becomes
@@
When considering the application to wildfire modeling in this paper, a simple analysis of order of magnitude of the terms in Eq. (89) shows that the total derivative of pressure as well as the viscous dissipation term can be neglected. If we compare the magnitude of these terms with the left-hand side of Eq. (89) we have
@@
for the characteristic scales of our problem, being the time scale related to the heating time from ambient temperature to the maximal temperature ($\Delta T\sim 600$ K). The term $\Delta p$ is negligible, as we consider constant atmospheric pressure.
@@
Analogously, we can rewrite the conservative form of the equation as
@@
Following Sec. 2.1, we assume that the medium can be represented by a two-phase porous flow, where the solid phase or fuel is denoted by $f$ and the gaseous phase, or air, is denoted by $a$. The fraction of fuel in the domain is given by $R_{f}$. We assume we have a perfect mixture and all variables are homogenized. All the material properties –e.g., density, specific heat– are defined in Sec. 2.1. By combining the conservation equations for enthalpy in a fluid and a solid, see [59], under the assumption of thermal equilibrium ($T_{f}=T_{a}=T$), we obtain a single-equation model as described in [51]:
@@
where $\rho h=\bar{\rho}_{f}h_{f}R_{f}+\bar{\rho}_{a}h_{a}(1-R_{f})$ is the enthalpy per unit volume of the mixture, $\mathbf{q}$ is a the total diffusion heat flux and $q_{s}$ represents all the heat sources (in both phases). Analogously, we can write
@@
Appendix B Validation of the numerical solvers – Mesh convergence analysis
In this section, we carry out a mesh convergence analysis of the model (using S2M), with and without moisture and wind. Fig. 21 shows the rate of spread depending on the number of mesh cells $N_{x}$ in the spatial discretization, compare Sec. 3.1. We consider low wind conditions, i.e. $w=0.2$ m/s and the default parameters from Tabs. 24. The initial condition in Eq. (74) is used. The domain length is 500 m.
Figure 21: Rate of spread depending on the number of cells. Top-left: $M=0.0$, $w=0.2$ and $t_{f}=2000$ s; top-right: $M=0.2$, $w=0.0$ and $t_{f}=2000$ s; bottom: $M=0.5$, $w=0.5$ and $t_{f}=2000$ s.
@@ Appendix B Validation of the numerical solvers – Mesh convergence analysis
Figure 22: Profiles of the solution depending on the number of cells. Top: $M=0.0$, $w=0.2$ and $t_{f}=2000$ s; middle: $M=0.2$, $w=0.0$ and $t_{f}=2000$ s; bottom: $M=0.5$, $w=0.5$ and $t_{f}=2000$ s.

### v2

@@ Acknowledgments
This work has been funded by the Spanish Ministry of Science, Innovation and Universities - Agencia Estatal de Investigación (10.13039/501100011033) and FEDER-EU under project-nr. PID2022-141051NA-I00. This work was partially funded by the Aragón Government, DGA (Spain), through the Fondo Europeo de Desarrollo Regional (FEDER) under project-nr. T32-23R. I. Özgen-Xian’s work is funded by the Tenure-Track-Programm by the German Federal Ministry of Education and Research. A. Navas-Montilla thanks Prof. César Dopazo for the insightful discussions on wildfire modelling, which offered valuable perspectives and inspired potential future developments.
@@ Appendix A Energy conservation equations in the model
In this section, the conservation of energy and the assumptions made to develop the model herein are presented. Refer to [8] for a complete description of these equations. For the sake of clarity, notation has been simplified and subscripts for the material properties for the fluid and solid cases have been omitted when analyzed independently.
@@
Now, by expanding the left-hand side of Eq. (91) and making use of the mass conservation equation—i.e. $\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{v})=0$, Eq. (91) is rewritten in its non-conservative form
@@
By expressing enthalpy differential in terms of pressure and temperature [8], Eq. (92) becomes
@@
where $c_{p}$ is the specific heat at constant pressure. Considering an ideal gas, the relation $\left(\frac{\partial\ln\rho}{\partial\ln T}\right)_{p}=-1$ holds and the previous equation becomes
@@
When considering the application to wildfire modeling in this paper, a simple analysis of order of magnitude of the terms in Eq. (94) shows that the total derivative of pressure as well as the viscous dissipation term can be neglected. Comparing the magnitude of these terms with the left-hand side of Eq. (94), we obtain
@@
for the characteristic scales of our problem, being the time scale related to the heating time from ambient temperature to the maximal temperature ($\Delta T\sim 600$ K). The term $\Delta p$ is negligible for a constant atmospheric pressure.
@@
Analogously, the conservative form of the equation becomes
@@
Following Sec. 2.1, the medium can be represented by a two-phase porous flow, where the solid phase or fuel is denoted by $f$ and the gaseous phase, or air, is denoted by $a$. The volume fraction of fuel in the domain is given by $R_{f}$. Assume having a perfect mixture and all variables are homogenized. All the material properties –e.g., density, specific heat– are defined in Sec. 2.1. By combining the conservation equations for enthalpy in a fluid and a solid, see [62], under the assumption of thermal equilibrium ($T_{f}=T_{a}=T$), we obtain a single-equation model as described in [54]:
@@
where $\rho h=\bar{\rho}_{f}h_{f}R_{f}+\bar{\rho}_{a}h_{a}(1-R_{f})$ is the enthalpy per unit volume of the mixture, $\mathbf{q}$ is a the total diffusion heat flux and $q_{s}$ represents all the heat sources (in both phases). Analogously,
@@
Appendix B Verification of the numerical solvers – Mesh convergence analysis
This section provides a mesh convergence analysis of the model (using S2M), with and without moisture and wind. Fig. 19 shows the rate of spread depending on the number of mesh cells $N_{x}$ in the spatial discretization, compare Sec. 3.1. Low wind conditions, i.e. $w=0.2$ m/s and the default parameters from Tabs. 24 are used. The initial condition in Eq. (79) is used. The domain length is 500 m.
(a) $M=0.0$, $w=0.2$
(b) $M=0.2$, $w=0.0$
(c) $M=0.5$, $w=0.5$
Figure 19: Rate of spread depending on the number of cells for different moisture and wind scenarios at $t_{f}=2000$ s. All scenarios show a convergence tendency for high number of cells.
@@ Appendix B Verification of the numerical solvers – Mesh convergence analysis
Figure 20: Profiles of the solution depending on the number of cells for different wind and moisture scenarios at a final time $t_{f}=2000$ s. Top: $M=0.0$, $w=0.2$; middle: $M=0.2$, $w=0.0$; bottom: $M=0.5$, $w=0.5$. The simulations show a convergence tendency for the traveling wave front and the maximal temperature.